OpenStudy (rainbow_rocks03):

The length of one side of a square is cm. What is the area of the square?

OpenStudy (rainbow_rocks03):

The length of one side of a square is cm. What is the area of the square? 

OpenStudy (rainbow_rocks03):

A. 7396 cm2 B. 43 cm2 C. 172 cm2 D. 86 cm2

OpenStudy (rainbow_rocks03):

@FlashBT3

OpenStudy (rainbow_rocks03):

@sleepyjess

OpenStudy (rainbow_rocks03):

@CallMeKiki

OpenStudy (callmekiki):

A = w * h

OpenStudy (rainbow_rocks03):

Do I do 86*4 ?

OpenStudy (callmekiki):

I can't see the sketch thingy.

OpenStudy (rainbow_rocks03):

What sketch thingy?

OpenStudy (callmekiki):

Oops. Wrong thing.

OpenStudy (callmekiki):

Okay.

OpenStudy (rainbow_rocks03):

So am I supposed to do 86*4 ? cause I asked you that and you dais something else.

OpenStudy (rainbow_rocks03):

@TheGaMer007 @sleepyjess @johnweldon1993 @FlashBT3 @Awolflover1

OpenStudy (rainbow_rocks03):

@CallMeKiki

OpenStudy (johnweldon1993):

\[\large Area = s^2\] Since the area is length x width and the side lengths are all equal so what is \[\large \sqrt{86}^2\]

OpenStudy (sleepyjess):

The area of a square is \(s^2\) we know s is \(\sqrt{86}\), so what is \(\sqrt{86}^2\)

OpenStudy (sleepyjess):

wow.....

OpenStudy (anonymous):

what happens when you multiply a square root by itself it becomes just what the number is making your answer D

OpenStudy (rainbow_rocks03):

86*2 ?

OpenStudy (johnweldon1993):

Not quite...square root of 86 times square root of 86...

OpenStudy (rainbow_rocks03):

ohhh

OpenStudy (rainbow_rocks03):

oh

OpenStudy (rainbow_rocks03):

PEOPLE PLZ HELP ME I HAVE TO DO 34 QUESTIONS AND THIS IS QUESTION #8 I CAN DO SOME QUESTIONS BUT THE ONES I DON"T UNDERSTAND I COME HERE ON THIS WEBSITE TO GET HELP.

OpenStudy (sleepyjess):

We are trying to help you...

OpenStudy (anonymous):

Hey I have stuff to do too, but I'm tring to do as much as I can for you. it's alright just calm down

OpenStudy (sleepyjess):

Let's use an example. If we are looking for the area of a square and one side of the square is \(\sqrt{4}\), we need to do \(\sqrt{4}^2\), and \(\sqrt{4}^2\) is 4

OpenStudy (sleepyjess):

\(\sqrt 4\) probably wasn't the best example... but hopefully that helps

OpenStudy (rainbow_rocks03):

never mind

OpenStudy (sleepyjess):

rainbow, is this a test?

OpenStudy (rainbow_rocks03):

thx for trying tho

OpenStudy (rainbow_rocks03):

and no it isn't

OpenStudy (anonymous):

\[\sqrt{86}*\sqrt{86}=?\]